One of the undersecretary of the Department of Labor and

Employment (DOLE) claims that the average salary of civil engineer

is Php18,000. A sample of 19 civil engineer’s salary has a mean of

Php17,350 and a standard deviation of Php1,230. Is there enough

evidence to reject the undersecretary’s claim at α = 0.01?


1
Expert's answer
2022-03-29T12:51:38-0400

Hypothesized Population Mean "\mu=18000"

Sample Standard Deviation s=1230

Sample Size n=19

Sample Mean "\bar{x}=17350"

Significance Level "\alpha=0.01"

Null and Alternative Hypotheses


The following null and alternative hypotheses need to be tested:

"H _\n0\n\u200b\n :\u03bc=18000"

"H _\n1\n\u200b\n :\u03bc\n\\not=\n18000"

This corresponds to two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.


Based on the information provided, the significance level is "\alpha=0.01"

"df=n\u22121=18"

degrees of fredom, and the critical value for two-tailed test is

"t_c=2.87844"

The rejection region for this left-tailed test is "R=\{t:|t|>2.87844\}"

The tt - statistic is computed as follows: "t=\frac{x-\mu}{s/\sqrt{n}}=\frac{17350-18000}{1230/\sqrt{19}}=-2.30348"

Since it is observed that

"|t|=2.30348<2.87844=t_c"

it is then concluded that the null hypothesis is not rejected.


Therefore, there is not enough evidence to claim that the population mean "\mu" is different than 18000, at the "\alpha=0.01"

significance level.




Using the P-value approach: The p-value for two-tailed, the significance level

"\alpha=0.01, df=18, t=-2.30348"

is p=0.033425 and since

"p=0.033425>0.01=\alpha"


it is concluded that the null hypothesis is not rejected.


Therefore, there is not enough evidence to claim that the population mean "\mu" is different than 18000, at the "\alpha=0.01"

significance level.